OpenStudy (anonymous):

Please help (: 1. Which is the equation of the given line in slope-intercept form? a. y=3x+2 b. y=3x-2 c. y=-3x+2 d. y=-1/3x+2 (THE LINE IS (-1,5 & 1,-1) 2. Which is the equation of the given line in standard form? A. y = 2 B. y = 2x C. x = 2 D. x = –2 (THE LINE IS 2,4 & 2,1)

OpenStudy (cwrw238):

wheres the line?

OpenStudy (anonymous):

slope intercept form is --> y=mx+b 'M' is the slope or \[\frac{ rise }{ run }\] 'B' is the y-intercept

OpenStudy (texaschic101):

lets do this together.....first you need to find the slope using the slope formula : slope(m) = (y2 - y1) / (x2 - x1) (-1,5) x1 = -1 and y1 = 5 (1,-1) x2 = 1 and y2 = -1 now we sub slope(m) = (-1 -5) / (1 - (-1) can you take it from here and find the slope ?

OpenStudy (anonymous):

yes thank you very much! :)

OpenStudy (texaschic101):

that is just the slope...that is not the equation in slope intercept form

OpenStudy (texaschic101):

do you still need help ?

OpenStudy (texaschic101):

hello ??

OpenStudy (anonymous):

#1 is C

OpenStudy (anonymous):

sorry my computer was frozen for a moment... thanks :) do you know #2?